The length of a rectangular land is 12m longer than the width. If the area is 864m2, find it's width.
Answer by checkley77(12844)

L=W+12
LW=864
(W+12)W=864
W^2+12W-864=0
(W-24)(W+36)=0
W-24=0
W=24 ANS. FOR THE WIDTH.
L=24+12=36 ANS. FOR THE LENGTH.
pROOF:
12*36=864
864=864